"An Analysis of the Study Time-Grade Association," published in Radical Pedagogy in 2002, looked at self-reported study time for the course, in hours per week, by 140 students enrolled in an Introductory Psychology course. Times had mean 3 hours and standard deviation 2.6 hours.

(a) The article reports that according to the National Survey of Student Engagement, the Carnegie Foundation benchmark is 6 hours of study per week for a course that meets three hours a week. Report the standardized sample mean time. (Round your answer to two decimal places.)

Used (x-mu)/(s/sqrt n)=(6-3)/(2.6/sqrt 140) and got 13.65 but was incorrect.

Also tried (6-3)/2.6 and was marked incorrect.
Answer #1

1- p(x>6)= p(z> (x-u)/sd)

= p( z>(6-3)/2.6)

= p( z>1.1538)

= 0.13 or 0.12 try both

plz comment below if you have query